1. fbox: xyz

2. mathbox: \sqrt{x}

3. setbox: width=17.50005pt, height=6.15079pt, depth=0.0pt

4. phantom: X Y, X Y

5. math symbol:(c.f fontmath.ltx)

(a) mathord: x M y,x∀y

(b) mathop: M_{x}^{y} ,∑_{x}^{y}

(c) mathbin: x M y, x + y

(d) mathrel: x M y, x = y

(e) mathopen,mathclose: M y N , (y)

(f) mathpunct: x M y, x, y

6. mathop: E_{y}^{x} z, E_{y}^{x} z, E_{y}^{x} z, E_{y}^{x} z

7. atop: {x \atop y}, \binom{x}{y}

8. genfrac(AmS): \frac{x}{y}, \binom{x}{y}, [\frac{x}{y}]

9. array: \begin{matrix}x \\ y \end{matrix}

10. oalign: X Y, X Y

11. ialign: X Y

12. overset,underset(AmS): E^{n} + E_{n}

13. stackrel: x =^{D} y

14. buildrel: x =^{D} x

15. kern: X YZ

16. Space corrections: 1m n, 2m n, 3m n, 4m n, 5m n, 6mn, 7mn, 8m n, 9mn, 0mn 17. mspace(AmS): m n, mn 18. Without italic correction(rir), with italic correction(ri r).

As a mid-sentence period(I. I), as an end-of-sentence punctuation(I. I). Inter-word space: without correction(etc. xαβ y z), with correction(etc. x α β y z)

19. hskip: x y

20. hspace: x y z

21. hfill:

A  B  C
A  B  C
A  B  C

22. ooalign: O / O / =

23. hbox,llap,rlap : XY Z, XY Z X YZ, X YZ

24. makebox: XY Z,X YZ

25. raise,lower: XY Z, X Y Z

26. raisebox: XYZ

27. smash: text

28. strut,mathstrut: \sqrt{x} + \sqrt{y}

29. rule: \sqrt{x}

30. vmargin: \sqrt{x}